DONAVAN, INC. Balance Sheet Assets Cash Inventory Buildings Total assets $12,000 15,000 40,000 $67,000 Liabilities and Stockholders Equity Liabilities Accounts payable Stockholders Equity $5,800 Common stock Retained earnings Total liabilities and stockholders equity $67,000 DONAVAN, INC. Income Statement $84,200 Revenues Cost of goods sold Salaries and wages expense 10,430 Net Income DONAVAN, INC. Retained Earnings Statement $17,000 Beginning retained earnings Add: Net income Less: Dividends Ending retained earnings 4,610 $32,000
Answer #1

Ending retained earnings=retained earnings =$32000(B).

Beginning retained earnings+Net income-Dividends=Ending retained earnings

17000+Net income-4610=32000

Net income=32000+4610-17000

=$19610.(D).(E).

Revenue-Cost of goods sold-Salaries and wages expenses=Net income

84200-Cost of goods sold-10430=19610

Cost of goods sold=84200-10430-19610

=$54160(C).

Accounts payable+Common stock+Retained earnings=Total liabilities and stockholders equity

5800+Common stock+32000=67000

Common stock=67000-32000-5800

=$29200.(A).